OpenStudy (anonymous):

Express the sum in the simplest form. 3x/ x+5 + 7/ x+5

OpenStudy (solomonzelman):

\(\Huge\color{black}{ \rm \frac{a}{b}+\frac{c}{b}=\frac{a+v}{b} }\)

OpenStudy (anonymous):

so add them together?

OpenStudy (anonymous):

\[\Huge\color{black}{ \rm \frac{a}{b}+\frac{c}{b}=\frac{a+c}{b} }\]:D

OpenStudy (anonymous):

denums are same, so add nums :-)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome